Please help! Thank-you!
Have a mixture of chocolate covered pretzels & macadamia nuts. The pretzels sell for $3/pd. and the nuts $8/pd. How many pds. of the nuts should be mixed with 50pds. of the pretzels to make a mixture that would sell for $6/pd?
A text messaging plan for a cell phone w/unlimited messages for $5/per month or a plan that charges $2 for the first 30 messages and then 10 cents for each additional message. How many messages (above the initial 30) would you need to send or recieve to make buying the unlimited messages plan a better deal?

Answer by stanbon(75887)

Let amount of nuts be "x".
Value of nuts is 8x dollars
Value of pretzels is 3(50) dollars
EQUATION:
Value of nuts + Value of pretzels = value of mixture
8x+150=6(x+50)
8x+150=6x+300
2x = 150
x= 75 lbs. of nuts should be added to the pretzels
Cheers,
Stan H.